Sirius The brightest star in the night sky is Sirius, which is at a distance of 8.30×1016 m, when we look at this star, how far back in time are we seeing it? Express your answer in years. Submit Answer Tries 0/15
Answer #1

The time lag will be

t=\frac{d}{c}=\frac{8.30\times10^{16}}{3\times10^8}=2.7667\times10^8$ s$=\frac{2.7668\times10^8}{365\times24\times3600}=8.773 $ yr$